OpenStudy (anonymous):

multiply and simplify by factoring: sqrt(10) * sqrt(14) =

OpenStudy (anonymous):

sqrt[140] just multiply the number inside

myininaya (myininaya):

\[\sqrt{2 \cdot 5} \sqrt{2 \cdot 7}=2\sqrt{35}\]

OpenStudy (anonymous):

2sqrt35

OpenStudy (anonymous):

2 srt(35)